OpenStudy (countrygirl1431):

@jim_thompson5910 could you possibly help me with 2 more?

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (countrygirl1431):

Write the equation of the line that passes through the given points. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. (2, 3) and (2, 5)

jimthompson5910 (jim_thompson5910):

what's the slope of the line?

jimthompson5910 (jim_thompson5910):

hint: slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (countrygirl1431):

5-3/2-2

jimthompson5910 (jim_thompson5910):

up top we have 5-3 = 2 down in the bottom we have 2-2 = 0 since we cannot divide by zero, the slope is undefined

jimthompson5910 (jim_thompson5910):

this is a vertical line. All vertical lines have undefined slopes

jimthompson5910 (jim_thompson5910):

the equation is simply x = 2 since EVERY point has an x coordinate of 2

OpenStudy (countrygirl1431):

ok thanks and the last one is Write the standard form of the line that passes through the given points. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. (4, 7) and (0, 7)

jimthompson5910 (jim_thompson5910):

what's the slope here?

OpenStudy (countrygirl1431):

id have to figure it just 1 sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (countrygirl1431):

7-7/0-4

jimthompson5910 (jim_thompson5910):

can you simplify that further?

OpenStudy (countrygirl1431):

no

jimthompson5910 (jim_thompson5910):

up top 7-7 turns into ??

OpenStudy (countrygirl1431):

0

jimthompson5910 (jim_thompson5910):

yep, and down below?

OpenStudy (countrygirl1431):

-4

jimthompson5910 (jim_thompson5910):

we have 0 over -4 0 divided by -4 = 0

jimthompson5910 (jim_thompson5910):

so the slope is m = 0

OpenStudy (countrygirl1431):

ok

jimthompson5910 (jim_thompson5910):

since the slope is 0, the equation is simply y = 7 every point on this line has a y coordinate of 7

OpenStudy (countrygirl1431):

so how would i put this in standard form

jimthompson5910 (jim_thompson5910):

it's already in standard form notice how `y = 7` is really `0x + 1y = 7`

OpenStudy (countrygirl1431):

oh ok i got it thanks for all your help

jimthompson5910 (jim_thompson5910):

you're welcome